# Degeneracy Enhancement of Neutron-Antineutron Oscillation in Neutron Star

### v1

@@ Abstract
We explore the fermion oscillation in a degenerate environment. The direct consequence is introducing a Pauli blocking factor $1-f_{i}$, where $f_{i}$ is the phase space distribution function, for each intermediate mass eigenstate during propagation. It is then much easier for a state with larger existing fraction or density to oscillate into other states with less degeneracy while the reversed process is not enhanced. This can significantly modify the oscillation behaviors. We apply this degenerate fermion oscillation to a concrete scenario of neutron-antineutron oscillation in neutron star. It turns out antineutrons receive a standing fraction to annihilate with the environmental neutrons. The subsequent neutron star heating can put an extremely stringent bound on the baryon number violating cross mass term between neutron and antineutron.
@@
Fermion Oscillation with Degeneracy – The formalism for fermion oscillation in vacuum is well established for neutrinos Pontecorvo:1957cp ; MNS ; P67 . The initial flavor state $|\alpha\rangle$ can be decomposed into a linear combination of mass eigenstates $|i\rangle$, $|\alpha\rangle=\sum_{i}U^{*}_{\alpha i}|i\rangle$ with $U_{\alpha i}$ being the mixing matrix element Kayser:2004wmw . With definite dispersion relation and hence propagating as plane wave, the mass eigenstate evolves independently, $|i(t,L)\rangle=e^{i\phi_{i}(t,L)}|i\rangle$ where $\phi_{i}(t,L)$ is the corresponding evolution phase over time interval $t$ and distance $L$. Correspondingly, the original flavor state $|\alpha(t,L)\rangle$ also evolves with time and distance. For measurement also in a flavor state $|\beta\rangle$, the amplitude $A_{\beta\alpha}\equiv\langle\beta|\alpha(t,L)\rangle$ is no longer simply $\langle\beta|\alpha\rangle=\delta_{\alpha\beta}$ but
$A_{\beta\alpha}=\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}.$ (1)
With different evolution phases $\phi_{i}\neq\phi_{j}$, their interference leads to oscillatory behavior in the transition probability $P_{\alpha\beta}\equiv|A_{\beta\alpha}|^{2}$. We have kept our notations general in order to cover neutrino and other fermion oscillations. For neutrino oscillation, the individual evolution phase is $\phi_{i}=m^{2}_{i}L/2E_{\nu}$ and the oscillation phase is $\Delta\phi_{ij}=\Delta m^{2}_{ij}L/2E_{\nu}$ with $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$. One may see that the neutrino oscillation phenomena is actually a quantum interference over macroscopic distance.
@@
In a concrete process, the fermion in propagation is created by a creation operator $a^{\dagger}_{\bm{p}i}$ and then propagate in the dense environment before being annihilated by $a_{\bm{p}i}$. The corresponding amplitude takes the form as $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle$. For simplicity, we have omitted other parts such as the production and detection vertices. Note that the creation and annihilation operators are always associated with a physical mass eigenstate that has definite dispersion relation. If the mass eigenstate propagate over distance, the evolution phase $e^{i\phi_{i}(t,L)}$ should also be taken into consideration.
The product of creation and annihilation operators is closely related to the particle momentum density, $\hat{n}_{\bm{p}i}\equiv a^{\dagger}_{\bm{p}i}a_{\bm{p}i}$. When applying to the environmental state $|\Omega\rangle$, the number density operator $\hat{n}_{\bm{p}i}$ should produce the particle momentum density, $\hat{n}_{\bm{p}i}|\Omega\rangle=n_{\bm{p}i}|\Omega\rangle$ with $n_{\bm{p}i}\equiv\int\mathrm{d}^{3}\bm{x}f_{i}(\bm{x},\bm{p})$ where $f_{i}(\bm{x},\bm{p})$ is the corresponding phase space distribution function. For a fermion, the creation and annihilation operators anticommute as, $\{a_{\bm{p}_{1}i},a_{\bm{p}_{2}i}^{\dagger}\}=(2\pi)^{3}\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$. Consequently, the product $a_{\bm{p}i}a^{\dagger}_{\bm{p}i}$ gives the Pauli blocking factor,
@@
where the $\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$ has been removed by momentum integration in the second quantized field operator, $\nu(x)\sim\int\mathrm{d}^{4}p(aue^{-ip\cdot x}+b^{\dagger}ve^{ip\cdot x})$.
@@
Taking the neutrino oscillation as a concrete example, the whole scattering amplitude of production, propagation, and detection takes the form as,
$\mathcal{M}$ $=\langle f,\Omega|(\bar{\ell}_{\beta}\gamma^{\mu}P_{L}W_{\mu}^{-}\nu_{\beta})(\bar{\nu}_{\alpha}\gamma^{\nu}P_{L}\ell_{\alpha}W_{\nu}^{+})|i,\Omega\rangle$ (3a)
@@
$\equiv\mathcal{M}_{d}\left[\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle\right]\mathcal{M}_{p},$ (3b)
where $W^{\pm}$ is the $W$ gauge boson and $\ell_{\alpha}$ the charged lepton of flavor $\alpha$. For convenience, the weak coupling constant has been omitted. The two neutrino spinors $\bar{u}$ and $u$ have been absorpted into the production and detection matrix elements $\mathcal{M}_{p}$ and $\mathcal{M}_{d}$, respectively. In between, the creation and annihilation operators sandwiched by the thermal environment state $|\Omega\rangle$ would give the Pauli exclusion factor, $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle=1-f_{i}$. Although illustrated with neutrino oscillation, this feature should apply generally to any fermion oscillation.
The oscillation amplitude in Eq. (1) then generalizes to
$A_{\beta\alpha}=\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}(1-f_{i}).$ (4)
The Pauli suppression factor $1-f_{i}$ represents the degeneracy effect brought by the environment. With an empty environment as vacuum ($f_{i}=0$), the amplitude reduces to the vacuum oscillation case as shown in Eq. (1). On the other hand, a fully degenerate Fermi gas $f_{i}=1$ does not allow oscillation to occupy the corresponding mass eigenstate.
With Pauli suppression factor, the amplitude in Eq. (4) can no longer reflect unitarity. Sticking to the usual definition of oscillation probability, $P_{\alpha\beta}\equiv|A_{\beta\alpha}|^{2}$, the total probability, $P_{\text{tot},\alpha}\equiv\sum_{\gamma}P_{\alpha\gamma}\neq 1$, is no longer one. It seems that the oscillating fermion would somehow disappear which is contradictory to the intuitive physical picture. However, oscillation in a degenerate environment does not introduce any mechanism for a particle to disappear. So unitarity has to be maintained.
To resolve this contradiction, one needs to resort to the more intuitive definition of oscillation probability. The original definition of probability or chance is actually the fraction of events. To be more exact, the oscillation probability $P_{\alpha\beta}$ should be the ratio between the number of $\beta$ flavor particles that can be detected after oscillation and the original number of $\alpha$ flavor ones without oscillation. It is then necessary to normalize the oscillation probability as $\mathcal{P}_{\alpha\beta}\equiv|A_{\alpha\beta}|^{2}/P_{\text{tot}}$ by hand. Note that the amplitude $A_{\alpha\beta}$ in Eq. (4) does not reduce to $\delta_{\alpha\beta}$ in the zero distance limit since the Pauli blocking is already in effect even if the oscillation does not really happen. In a degenerate environment, an $\alpha$ flavor state from production cannot contain the existing mass eigenstate.
@@
in the $(n,\bar{n})$ basis. The diagonal elements $H_{11}$ and $H_{22}$ are the neutron and antineutron energies Phillips:2014fgb . The cross mass $\delta m$ will induce the $n$-$\bar{n}$ mixing, $\tan 2\theta=2\delta m/(H_{22}-H_{11})$,
$U=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix},\quad\begin{pmatrix}n_{1}\\ n_{2}\end{pmatrix}=U\begin{pmatrix}n\\ \bar{n}\end{pmatrix},$ (6)
where $n_{1}$ and $n_{2}$ are the two mass eigen-fields. Typically, the mixing angle $\theta$ is highly suppressed by the tiny $\delta m$. For simplicity, one may denote the sine and cosine of the mixing angle as $(s,c)\equiv(\sin\theta,\cos\theta)$. The energy difference between the mass eigenstates is simply denoted as $\Delta E\equiv\sqrt{(H_{11}-H_{22})^{2}+4\delta m^{2}}$ Phillips:2014fgb . So the $n\bar{n}$ oscillation probabilities are
$P_{n\bar{n}}$ $=c^{2}s^{2}(1-f_{1})^{2}+s^{2}c^{2}(1-f_{2})^{2}$
$-2c^{2}s^{2}(1-f_{1})(1-f_{2})\cos(\Delta Et),$ (7a)
$P_{nn}$ $=c^{4}(1-f_{1})^{2}+s^{4}(1-f_{2})^{2}$
$+2c^{2}s^{2}(1-f_{1})(1-f_{2})\cos(\Delta Et),$ (7b)
where $f_{1}$ and $f_{2}$ are the phase space distribution functions of the two mass eigenstates. Note that the time dependence only appears in the $\cos(\Delta Et)$ term and the norminal total probability $P_{\rm tot}\equiv P_{nn}+P_{n\bar{n}}$ is indeed time-independent which is a necessary feature for our argument of probability normalization/definition above. The two probabilities reduce to the non-degenerate case with both $f_{1}$ and $f_{2}$ approaching 00.
The normalized oscillation probability should take the form as,
$\mathcal{P}_{n\bar{n}}$ $=\frac{c^{2}s^{2}(1-f_{1})^{2}+c^{2}s^{2}(1-f_{2})^{2}}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}}$
@@
with $\mathcal{P}_{nn}\equiv P_{nn}/P_{\rm tot}$ being omitted for simplicity but can be readily reconstructed. The total probability $P_{\rm tot}$ in the denominator reduces to 1 in the non-degenerate limit. It is interesting to see that the phase space distribution function appears in the denominator as $1-f_{i}$. In a dense environment with degeneracy, the phase space distribution function $f_{i}$ can approach 1 below the Fermi surface. It is then possible for the oscillation probabilities to receive enhancement from degeneracy.
@@
With a tiny mixing angle $\theta$ Phillips:2014fgb , $s\ll 1$, the mass eigenstate $n_{1}$ is approximately the neutron state $n$ that dominates the matter inside NS. Consequently, both $1-f_{1}$ and $f_{2}$ are negligibly small. Between the two tiny terms $s$ and $1-f_{1}$ in the denominator, the case $s\gg 1-f_{1}$ would lead to significant transition rate $\mathcal{P}_{n\bar{n}}\approx c^{2}$ which means degenerate neutrons would instantly transit to antineutrons. The situation $s\sim 1-f_{1}$ has also the same conclusion. Since neutron stars are readily observed and hence should be stable, only the remaining case $s\ll 1-f_{1}$ can happen. This already indicates that the degenerate $n\bar{n}$ oscillation in NS can give a very much enhanced constraint.
To be more exact, in the limit of $s\ll 1-f_{1}\ll 1$ and $f_{2}\ll 1$, the transition rate from neutron to antineutron becomes
$\mathcal{P}_{n\bar{n}}\approx\frac{s^{2}}{(1-f_{1})^{2}}\equiv s^{2}_{\rm eff},\qquad\mathcal{P}_{nn}\approx c^{2}.$ (9)
@@
It is interesting to observe that the transition rate $\mathcal{P}_{n\bar{n}}$ is enhanced by degeneracy. This is consistent with the intuitive picture that it is much easier for neutrons in a degenerate neutron environment to transit to other particles without Pauli blocking. The effective mixing $s^{2}_{\rm eff}$ should be much larger than the vacuum counterpart $s^{2}$. Switching neutron with antineutron and hence $f_{1}$ with $f_{2}$, the reversed transition probability $\mathcal{P}_{\bar{n}n}\approx s^{2}$ is not enhanced. The bidirectional oscillation has imbalanced rates to derive the overall transition from the degenerate neutrons to the non-degenerate antineutrons.
@@
Standing Fraction of Antineutron in Neutron Star – Another important feature is that the time dependence of the $n\bar{n}$ oscillation probability $\mathcal{P}_{n\bar{n}}$, which is $2c^{2}s^{2}\sin^{2}(\Delta Et/2)$ for the vacuum case, is now suppressed by extra $1-f_{1}$. In other words, the transition happens instantaneously. In an extremely degenerate environment such as NS, neutrons will transit into antineutrons so rapidly that the ratio between their distribution functions $f_{\bar{n}}/f_{n}$ stays the same,
$R(\bm{p})\equiv\frac{f_{\bar{n}}}{f_{n}}=\frac{\mathcal{P}_{n\bar{n}}}{\mathcal{P}_{nn}}\approx\frac{\tan^{2}\theta}{[1-f_{1}(\bm{p})]^{2}},$ (10)
@@
which actually reduces to the effective mixing $s^{2}_{\rm eff}$. With momentum dependence inherited from $f_{1}(\bm{p})$, $s^{2}_{\rm eff}$ does not have a fixed value. Even directly taking the zero time limit of Eq. (8), $\mathcal{P}_{n\bar{n}}(t=0)\approx s^{2}(f_{1}-f_{2})^{2}/(1-f_{1})^{2}$ where the approximation only applies for the denominator, Eq. (10) can be readily reproduced with $(f_{1}-f_{2})^{2}\approx 1$. In other words, the production process is already affected by the degenerate environment as pointed out earlier. The neutron transition to antineutron can instantly happen and remain the same antienutron fraction over time.
@@
Being in thermal equilibrium as maintained inside NS, neutrons follow the Fermi-Dirac distribution, $f_{1}(\bm{p})=1/[e^{(\varepsilon_{n}(\bm{p})-\mu)/T}+1]$ where $\varepsilon_{n}(\bm{p})=\sqrt{\bm{p}^{2}+m^{2}}$ is the neutron energy with momentum $\bm{p}$ and mass $m$. With a very low internal temperature $T\sim 7$ keV at the age of $10^{5}$ years Potekhin:2020ttj , the corresponding neutron Fermi momentum $k_{F}=(3\pi^{2}n_{n})^{1/3}$ and chemical potential $\mu_{n}=\sqrt{k_{F}^{2}+m^{2}}$ can be readily derived from the neutron density $n_{n}$ assuming full degeneracy. We consider a typical NS with mass $M=1.4M_{\odot}$ (solar mass $M_{\odot}$) and radius $R_{n}=10$ km Lattimer:2000nx . To match the corresponding high density $n_{n}\approx 4\times 10^{38}$ cm-3 Potekhin:2015qsa , which is also assumed to be constant inside NS, the difference of chemical potential $\mu$ from the neutron mass $m$, $\mu-m\approx 101.6$ MeV, is much larger than temperature $T$ Brandes:2023bob . It is fully consistent to say that the distribution function is very close to 1 below the Fermi surface, $1-f_{1}(\bm{p})\approx e^{[\varepsilon_{n}(\bm{p})-\mu]/T}\ll 1$.
@@
For comparison, the non-degenerate approach gives an average antineutron fraction $R^{\prime}=2c^{2}s^{2}$ by neglecting the oscillating part Buccella:1987pv . Between the degenerate and non-degenerate cases, the difference in the antineutron fraction $R/R^{\prime}\equiv n_{\bar{n}}/n^{\prime}_{\bar{n}}$ can reach $\mathcal{O}(10^{3})$ orders with $(\mu-m)/T\sim 10^{4}$ which is a huge enhancement $e^{(\mu-m)/T}$ due to degeneracy. As shown in Fig. 1, the neutron degeneracy parameter $(\mu-m)/T$ decreases with radius but increases with time according to the NSCool simulation Page:2004fy ; NSCool program .
Figure 1: The degeneracy parameter $(\mu-m)/T$ of neutron matter inside NS as function of radius and cooling time.
Neutron Burning and Neutron Star Heating – A standing fraction of antineutron in the dense neutron environment as indicated by Eq. (10) would lead to constant burning of neutron matter by neutron-antineutron annihilation. In a single annihilation, the total nucleon number $N=N_{n}+N_{\bar{n}}$ decreases by two, $\mathrm{d}N=-2\langle\sigma v\rangle n_{n}n_{\bar{n}}\mathrm{d}tdV$, where $\langle\sigma v\rangle$ is the thermally averaged cross section $\sigma$ times the relative velocity $v$. Replacing the antineutron number density $n_{\bar{n}}=Rn_{n}$ gives,

### v2

# Degeneracy Enhancement of Neutron Oscillation in Neutron Star

@@ Abstract
We explore the fermion oscillation in a degenerate environment. The direct consequence is introducing a Pauli blocking factor $1-f_{i}$, where $f_{i}$ is the phase space distribution function, for each intermediate mass eigenstate during propagation. It is then much easier for a state with larger existing fraction or density to oscillate into other states with less degeneracy while the reversed process is not enhanced. This can significantly modify the oscillation behaviors. We apply this degenerate fermion oscillation to a concrete scenario of neutron-antineutron oscillation in neutron star. It turns out antineutrons receive a standing fraction to annihilate with the environmental neutrons. The subsequent neutron star heating can put an extremely stringent bound on the baryon number violating cross mass term between neutron and antineutron. The neutron-mirror-neutron oscillation in neutron star with cooling effect should have similar results.
@@
Fermion Oscillation with Degeneracy – The formalism for fermion oscillation in vacuum is well established Pontecorvo:1957cp ; MNS ; P67 . The initial flavor state $|\alpha\rangle$ can be decomposed into a linear combination of mass eigenstates $|i\rangle$, $|\alpha\rangle=\sum_{i}U^{*}_{\alpha i}|i\rangle$ with $U_{\alpha i}$ being the mixing matrix element Kayser:2004wmw . With definite dispersion relation and hence propagating as plane wave, the mass eigenstate evolves independently, $|i(t,L)\rangle=e^{i\phi_{i}(t,L)}|i\rangle$ where $\phi_{i}(t,L)$ is the corresponding evolution phase over time interval $t$ and distance $L$. Correspondingly, the original flavor state $|\alpha(t,L)\rangle$ also evolves with time and distance. For measurement also in a flavor state $|\beta\rangle$, the transition probability $P_{\alpha\beta}$ is no longer simply $\delta_{\alpha\beta}$ but
$P_{\beta\alpha}=\left|\sum_{i}U_{\beta i}e^{i\phi_{i}(t,L)}U^{*}_{\alpha i}\right|^{2}.$ (1)
With different evolution phases $\phi_{i}\neq\phi_{j}$, their interference leads to oscillatory behavior in the transition probability $P_{\alpha\beta}$. For neutrino oscillation, the individual evolution phase is $\phi_{i}\equiv m^{2}_{i}L/2E_{\nu}$ and the oscillation phase is $\Delta\phi_{ij}=\Delta m^{2}_{ij}L/2E_{\nu}$ with $\Delta m^{2}_{ij}\equiv m^{2}_{i}-m^{2}_{j}$. Note that our notations in Eq. (1) are kept general to cover both neutrino and other fermion oscillations.
@@
The fermion in propagation is created by a creation operator $a^{\dagger}_{\bm{p}i}$ and then propagate in the dense environment before being annihilated by $a_{\bm{p}i}$. The corresponding amplitude takes the form as $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle$. For simplicity, we have omitted other parts such as the production and detection vertices. Note that the creation and annihilation operators are always associated with a physical mass eigenstate that has definite dispersion relation. If the mass eigenstate propagate over distance, the evolution phase $e^{i\phi_{i}(t,L)}$ should also be taken into consideration.
The product of creation and annihilation operators is closely related to the particle momentum density, $\hat{n}_{\bm{p}i}\equiv a^{\dagger}_{\bm{p}i}a_{\bm{p}i}$. When applying to the environmental state $|\Omega\rangle$, the number density operator $\hat{n}_{\bm{p}i}$ should produce the particle momentum density, $\hat{n}_{\bm{p}i}|\Omega\rangle=n_{\bm{p}i}|\Omega\rangle$ with $n_{\bm{p}i}\equiv\int\mathrm{d}^{3}\bm{x}f_{i}(\bm{x},\bm{p})$ where $f_{i}(\bm{x},\bm{p})$ is the corresponding phase space distribution function. For a fermion, the creation and annihilation operators anticommute as, $\{a_{\bm{p}_{1}i},a_{\bm{p}_{2}i}^{\dagger}\}=(2\pi)^{3}\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$  [^1] During revision, we notice the same $\delta^{(3)}$ function has also been obtained in Ghosh:2022nzo ; Barbosa:2024tty .. Then the product $a_{\bm{p}i}a^{\dagger}_{\bm{p}i}$ gives the Pauli blocking factor Barbosa:2024tty ,
@@
where the $\delta^{(3)}(\bm{p}_{2}-\bm{p}_{1})$ has been removed by momentum integration in the second quantized fermion field operator, $\psi(x)\sim\int\mathrm{d}^{4}p(aue^{-ip\cdot x}+b^{\dagger}ve^{ip\cdot x})$.
@@
The scattering amplitude for the whole process composed of production, propagation, and detection takes the form as,
@@
$\mathcal{M}_{\beta\alpha}$ $\equiv\mathcal{M}_{d}\left[\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle\right]\mathcal{M}_{p},$ (3)
The fermion oscillation process is a very special type with the intermediate particle being on shell and the particle momentum does not change. These intermediate particles remain in coherence and evolve according to definite dispersion relation. The creation and annihilation operators from the intermediate fields applies on the thermal vacuum $|\Omega\rangle$ to produce the Pauli exclusion factor, $\langle\Omega|a_{\bm{p}i}a^{\dagger}_{\bm{p}i}|\Omega\rangle=1-f_{i}$. The two on-shell spinors $\bar{u}$ and $u$ have been absorpted into the production and detection matrix elements $\mathcal{M}_{p}$ and $\mathcal{M}_{d}$, respectively.
Attaching a Pauli blocking factor $1-f_{i}$ with an on-shell intermediate particle is fully consistent with the thermal quantum field theory (QFT) Lundberg:2020mwu . The fermion propagator in an ensemble of particles, $S_{F}(p)\propto i/(p^{2}-m^{2}+i\epsilon)-2\pi\delta(p^{2}-m^{2})\Theta(p^{0})f(\bm{p})$ Ghosh:2022nzo ; Notzold:1987ik would reduce to
$S_{F}$ $\propto\int\frac{d^{3}\bm{p}}{(2\pi)^{3}}\frac{1}{2E_{\bm{p}}}e^{-ip\cdot(x-y)}[1-f(\bm{p})]\left(\not{p}+m_{\nu}\right),$ (4)
after Fourier transformation $\int d^{4}p/(2\pi)^{4}$. The first term in the square bracket comes from the ordinary propagator $i/(p^{2}-m^{2}+i\epsilon)$ by extracting the physical pole while the $f(\bm{p})$ term from the thermal contribution after integrating away the on-shell $\delta(p^{2}-m^{2})$ function.
Treating the $1-f$ factor as field strength renormalization, our result can be readily obtained with the quantum kinetics theory Sigl:1993ctk ; Volpe:2023met . From the mass matrix diagonalization, the flavor fields $\psi_{\alpha}$ connect with the mass eigen-fields $\psi_{i}$ via a unitarity mixing matrix $U$, $\psi_{\alpha}=\sum_{i}U_{\alpha i}\psi_{i}$. By field redefinition, $\psi_{i}=\sqrt{Z_{i}}\psi^{\prime}_{i}$ with $Z_{i}\equiv 1-f_{i}$, the canonical field $\psi^{\prime}_{i}$ develops evolution phase $e^{i\phi_{i}(t,L)}$. Then the scattering matrix element should contain a factor of $U_{\beta i}(1-f_{i})e^{i\phi_{i}}U^{*}_{\alpha i}$ in the flavor basis, which is consistent with Eq. (3).
In the QFT language, the detected particle number of flavor $\beta$ is proportional to $N_{\beta}\propto|\mathcal{M}_{\beta\alpha}|^{2}$. To be more specific, the total particle number without propagation ($x=y$) is proportional to $N\propto\sum_{\beta}|\mathcal{M}_{\beta\alpha}(x=y)|^{2}$. After propagating over a nonzero space-time interval $x-y$, the detected particle number of flavor $\beta$ is proportional to $N_{\beta}(x-y)\propto|\mathcal{M}_{\beta\alpha}(x-y)|^{2}$. In other words, a fraction of $N_{\beta}(x-y)/N$ has oscillated to flavor $\beta$ which gives the probability for a neutron to be probed as an antineutron,
$P_{\alpha\beta}\equiv\frac{\left|\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}(1-f_{i})\right|^{2}}{\sum_{\beta}\left|\sum_{i}U_{\beta i}U^{*}_{\alpha i}e^{ip_{i}\cdot(x-y)}(1-f_{i})\right|^{2}}.$ (5)
With an empty environment ($f_{i}=0$), the transition probability reduces to the vacuum case as shown in Eq. (1). The same formalism of degenerate oscillation should also apply for bosons with the Pauli blocking factor $1-f_{i}$ replaced by the Bose enhancement factor $1+f_{i}$.
@@
in the $(n,\bar{n})$ basis. The diagonal elements $H_{11}$ and $H_{22}$ are the neutron and antineutron energies Phillips:2014fgb . The Majorana mass $\delta m$ that violates the baryon number by 2 will induce the $n$-$\bar{n}$ mixing, $\tan 2\theta=2\delta m/(H_{22}-H_{11})$,
$\begin{pmatrix}n_{1}\\ n_{2}\end{pmatrix}=U\begin{pmatrix}n\\ \bar{n}\end{pmatrix},\quad U=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix},$ (7)
where $n_{1}$ and $n_{2}$ are the two mass eigen-fields. Typically, the mixing angle $\theta$ is highly suppressed by the tiny $\delta m$. For simplicity, one may denote the sine and cosine of the mixing angle as $(s,c)\equiv(\sin\theta,\cos\theta)$. The energy difference between the mass eigenstates is simply denoted as $\Delta E\equiv\sqrt{(H_{11}-H_{22})^{2}+4\delta m^{2}}$ Phillips:2014fgb .
The oscillation probability takes the form as,
$P_{n\bar{n}}$ $=\frac{c^{2}s^{2}(1-f_{1})^{2}+c^{2}s^{2}(1-f_{2})^{2}}{c^{2}(1-f_{1})^{2}+s^{2}(1-f_{2})^{2}}$
@@
where $f_{1}$ and $f_{2}$ are the phase space distribution functions of the two mass eigenstates. Note that the phase space distribution function appears in the denominator as $1-f_{i}$. In a dense environment with degeneracy, the phase space distribution function $f_{i}$ can approach 1 below the Fermi surface. It is then possible for the oscillation probabilities to receive enhancement from degeneracy.
@@
With a tiny mixing angle $\theta$ Phillips:2014fgb , $s\ll 1$, the mass eigenstate $n_{1}$ is approximately the neutron state $n$ that dominates the matter inside NS. Consequently, both $1-f_{1}$ and $f_{2}$ are negligibly small. Between the two tiny terms $s$ and $1-f_{1}$ in the denominator, the case $s\gg 1-f_{1}$ would lead to a large $P_{n\bar{n}}\approx c^{2}$ which means degenerate neutrons would instantly transit to antineutrons. The situation $s\sim 1-f_{1}$ has also the same conclusion. Since neutron stars are readily observed and hence should be stable, only the remaining case $s\ll 1-f_{1}$ can happen. This already indicates that the degenerate $n\bar{n}$ oscillation in NS can give a very much enhanced constraint.
Figure 1: Illustration of the degenerate $n$-$\bar{n}$ oscillation.
In the limit of $s\ll 1-f_{1}\ll 1$ and $f_{2}\ll 1$, the transition rate from neutron to antineutron becomes
$P_{n\bar{n}}\approx\frac{s^{2}}{(1-f_{1})^{2}}\equiv s^{2}_{\rm eff},\qquad P_{nn}\approx c^{2}.$ (9)
@@
As shown in Fig. 1, although the neutron state ($n$) is mainly composed of $n_{1}$, its coefficient $\cos\theta$ receives extra suppression $1-f_{1}$ due to Pauli blocking. So the relative fraction of $n_{2}$ and subsequently the chance for the antineutron ($\bar{n}$) to appear is enhanced. It is much easier for neutrons in a degenerate neutron environment to transit to antineutrons without Pauli blocking. Switching neutron with antineutron and hence $f_{1}$ with $f_{2}$, the reversed transition probability $P_{\bar{n}n}\approx s^{2}$ is not enhanced. The bidirectional oscillation has imbalanced rates to drive the overall transition from the degenerate neutrons to the non-degenerate antineutrons.
@@
Standing Fraction of Antineutron in Neutron Star – Another important feature is that the time dependence of the $n\bar{n}$ oscillation probability $P_{n\bar{n}}$, which is $2c^{2}s^{2}\sin^{2}(\Delta Et/2)$ for the vacuum case, is now suppressed by extra $1-f_{1}$. In other words, the transition happens instantaneously. Further more, the environment is full of neutrons such that no particle can freely propagate for sizable distance. It is unavoidable to consider the zero time/distance limit.
In an extremely degenerate environment, neutrons transit into antineutrons so rapidly that the ratio between their distribution functions $f_{\bar{n}}/f_{n}$ stays the same,
$R(\bm{p})\equiv\frac{f_{\bar{n}}}{f_{n}}=\frac{P_{n\bar{n}}}{P_{nn}}\approx\frac{\tan^{2}\theta}{[1-f_{1}(\bm{p})]^{2}},$ (10)
@@
which actually reduces to the effective mixing $s^{2}_{\rm eff}$. With momentum dependence inherited from $f_{1}(\bm{p})$, $s^{2}_{\rm eff}$ does not have a fixed value. Even directly taking the zero time limit of Eq. (8), $P_{n\bar{n}}(t=0)\approx s^{2}(f_{1}-f_{2})^{2}/(1-f_{1})^{2}$ where the approximation only applies for the denominator, Eq. (10) can be readily reproduced with $(f_{1}-f_{2})^{2}\approx 1$. In other words, the production process is already affected by the degenerate environment as pointed out earlier. The neutron transition to antineutron can instantly happen and remain the same antineutron fraction over time.
@@
Being in thermal equilibrium as maintained inside NS, neutrons follow the Fermi-Dirac distribution, $f_{1}(\bm{p})=1/[e^{(\varepsilon_{n}(\bm{p})-\mu)/T}+1]$ where $\varepsilon_{n}(\bm{p})=\sqrt{\bm{p}^{2}+m^{2}}$ is the neutron energy with momentum $\bm{p}$ and mass $m$. With a very low internal temperature $T\sim 7$ keV at the age of $10^{5}$ years Potekhin:2020ttj , the corresponding neutron Fermi momentum $k_{F}=(3\pi^{2}n_{n})^{1/3}$ and chemical potential $\mu=\sqrt{k_{F}^{2}+m^{2}}$ can be readily derived from the neutron density $n_{n}$ assuming full degeneracy. We consider a typical NS with mass $M=1.4M_{\odot}$ (solar mass $M_{\odot}$) and radius $R_{n}=10$ km Lattimer:2000nx . To match the corresponding high density $n_{n}\approx 4\times 10^{38}$ cm-3 Potekhin:2015qsa , which is also assumed to be constant inside NS, the difference of chemical potential $\mu$ from the neutron mass $m$, $\mu-m\approx 101.6$ MeV, is much larger than temperature $T$ Brandes:2023bob . It is fully consistent to say that the distribution function is very close to 1 below the Fermi surface, $1-f_{1}(\bm{p})\approx e^{[\varepsilon_{n}(\bm{p})-\mu]/T}\ll 1$. Although neutrons can form superfluid, it can only happen around the Fermi surface with a gap $\sim 1$ MeV Haskell:2017lkl much smaller than the chemical potential $\mu-m$ and Fermi momentum. In addition, $f_{1}$ deviates from 1 in this region and hence the $n\bar{n}$ oscillation would not experience degeneracy enhancement.
@@
For comparison, the non-degenerate approach gives an average antineutron fraction $R^{\prime}=2c^{2}s^{2}$ by neglecting the oscillating part Buccella:1987pv . Between the degenerate and non-degenerate cases, the difference in the antineutron fraction $R/R^{\prime}\equiv n_{\bar{n}}/n^{\prime}_{\bar{n}}$ can reach $\mathcal{O}(10^{3})$ orders with $(\mu-m)/T\sim 10^{4}$ which is a huge enhancement $e^{(\mu-m)/T}$ due to degeneracy. As shown in Fig. 2, the neutron degeneracy parameter $(\mu-m)/T$ decreases with radius but increases with time according to NSCool Page:2004fy ; NSCool program .
Figure 2: The degeneracy parameter $(\mu-m)/T$ of neutron matter inside NS without considering $n\bar{n}$ annihilation
as function of radius and cooling time.
Neutron Burning and Neutron Star Heating – A standing fraction of antineutron in the dense neutron environment as indicated by Eq. (10) would lead to constant burning of neutron matter by neutron-antineutron annihilation. In a single annihilation, the total nucleon number $N=N_{n}+N_{\bar{n}}$ decreases by two, $\mathrm{d}N=-2\langle\sigma v\rangle n_{n}n_{\bar{n}}\mathrm{d}t\mathrm{d}V$, where $\langle\sigma v\rangle$ is the thermally averaged cross section $\sigma$ times the relative velocity $v$. Replacing the antineutron number density $n_{\bar{n}}=Rn_{n}$ gives,